A bottling machine can be regulated so that it discharges an average of µ ounces per bottle. It has been observed that the amount of fill dispensed by the machine is normally distributed with σ = 1 ounce. A sample of n = 9 filled bottles is randomly selected from the output of the machine on a given day (all bottled with the same machine setting) and the ounces of fill measuered for each. 1 i) Find the probability that the sample mean will be within 0.3 ounce of the true mean for that particular setting. [4] ii) How many observations should be included in the sample if we wish X¯ to be within 0.3 ounce of µ with probability 0.95? [4] iii) Suppose we plan to select a random sample of ten bottles and measure the amount of fill in each bottle. If these ten observations are used to calculate S 2 , find numbers b1 and b2 such that P(b1 ≤ S 2 ≤ b2) = 0.90


1
Expert's answer
2022-01-31T16:26:42-0500

"a)"

Let "X_1,X_2,X_3,..X_9" denote the ounces of fill to be observed. Then, "X_i\sim N(\mu,1),\space 1\lt i\lt9". The sample mean "\bar {X}" is also normally distributed with mean "\mu" and variance "{\sigma^2\over n}". That is, "\bar X\sim N(\mu,{1\over 9})" as "n=9."

We need to determine,

"p(|\bar X-\mu|\leq0.3)=p(-0.3\leq\bar X-\mu\leq0.3)". this can be standardized as follows.

"p(-0.3\leq\bar X-\mu\leq0.3)=p({-0.3\over{1\over\sqrt{9}}}\leq{\bar X-\mu\over {1\over\sqrt{9}}}\leq{0.3\over{1\over\sqrt{9}}})\equiv p({-0.9\lt Z\lt0.9})"

This probability can be written as,

"p({-0.9\lt Z\lt0.9})=\phi(0.9)-\phi(-0.9)=0.8159-0.1841=0.6318"

Therefore, the probability that the sample mean will be within 0.3 ounce of the true mean is 0.6318.


"b)"

Here, we need the sample size "n" such that, "p(|\bar X-\mu|\leq0.3)=0.95".This probability can be written as, "p(-0.3\leq\bar X-\mu\leq0.3)=0.95..........(1)"

Dividing each term of the inequality in equation (1) by the standard deviation "\sigma_{\bar x}={1\over\sqrt{n}}", we have,

"p(-0.3\sqrt{n}\leq{\bar X-\mu\over {1\over\sqrt{n}}}\leq0.3\sqrt{n})=0.95" but "{\bar x-\mu\over{1\over\sqrt{n}}}=Z". Thus, "p(-0.3\sqrt{n}\leq{\bar X-\mu\over {1\over\sqrt{n}}}\leq0.3\sqrt{n})=p(-0.3\sqrt{n}\leq Z\leq0.3\sqrt{n})=0.95" . Using the standard normal tables, the value at "\alpha=0.05" is "Z_{0.05\over2}=Z_{0.025}=1.96". Therefore, "p(-0.3\sqrt{n}\leq Z\leq0.3\sqrt{n})=p(-1.96\leq Z\leq1.96)=0.95\implies 0.3\sqrt{n}=1.96"

Thus,

"0.3\sqrt{n}=1.96\implies n=({1.96\over0.3})^2=42.6844\approx 43."

Therefore, the number of observations which should be included in the sample if we wish "\bar X" to be within 0.3 ounce of "\mu" with probability 0.95 is "n=43."


"c)"

"p(b_1 \leq s^ 2 \leq b_2) = 0.90"

"n=10"

We standardize "s^2" to a "\chi^2" random variable as follows.

"p(b_1 \leq s ^2 \leq b_2)=p({(n-1)b_1\over \sigma^2}\leq {(n-1)s^2\over \sigma^2}\leq{ (n-1)b_2\over\sigma^2}) = 0.90"

Since "\sigma^2=1," it follows that "{(n-1)s^2\over\sigma^2}=(n-1)s^2" has a "\chi^2" distribution with "(n-1)=10-1=9" degrees of freedom.

Let "{(n-1)b_1\over \sigma^2}=(n-1)b_1" be "x_1" and "{(n-1)b_2\over \sigma^2}=(n-1)b_1" be "x_2" . We rewrite our probability as, "p(x_1\leq (n-1)s^2\leq x_2)=0.90"

To solve for values of "x_1" and "x_2", we proceed as below.

To determine the value of "x_1," we find a value that cuts off an area of 0.05 on the lower tail of the "\chi^2" distribution with 9 degrees of freedom. From the tables, this value is, "\chi^2_{0.05,9}=3.3251". Thus "x_1=3.3251". Since "x_1=3.3251\implies (n-1)b_1=3.3251\implies 9b_1=3.3251\implies b_1=0.3695"


Also, to determine the value of "x_2," we find a value that cuts off an area of 0.05 on the upper tail of the "\chi^2" distribution with 9 degrees of freedom. From the tables, this value is, "\chi^2_{0.05,9}=16.9190". Thus "x_2=16.9190". Since "x_2=16.9190\implies (n-1)b_2=16.9190\implies 9b_2=3.3251\implies b_2=1.8799"

Therefore, the values of "b_1" and "b_2" such that "P(b_1 \leq s^ 2\leq b_2) = 0.90" are 0.3695 and 1.8799 respectively.